# Towards Cooperative Maneuver Planning in Mixed Traffic at Urban Intersections

### v1

@@ V-C Reward Function Engineering
with describing the set of reward components and the corresponding weights. The only positive reward component and thus the main driver for learning a nontrivial solution is the velocity reward, which is proportional to the average driving speed clipped to the lane speed limit. A trivial solution that consequently avoids any collision is to stop all vehicles in the scene. This is not a valid solution for cooperative maneuver planning and thus prevented through the idle penalty. In mixed traffic, the training tends to converge to a very conservative behavior, which manifests in all CAVs being stopped far away from the intersection to let HDVs pass [3]. This issue is addressed by the reluctance penalty, which is triggered for CAVs that drive very slowly without any lead vehicle. In order to facilitate learning of reasonable safety distances, the proximity reward penalizes situations in which CAVs get dangerously close. Ultimately, two vehicles colliding triggers the collision penalty and ends the training episode to avoid further positive rewards from being accumulated. The reward weights used throughout this study, at a policy frequency of , are given in Table II. Details on the definition and tuning of the various reward components can be found in [20, 3].
@@
V-D Derivation of Cooperative Maneuver
The maneuver representation introduced in Sec. III requires the cooperative maneuver to be specified over a sufficient horizon to facilitate local trajectory planning on the CAVs. This requirement for advance planning contrasts with the iterative nature of the RL policy, which requires immediate state feedback. In the present work, we build upon our approach from [26] to bridge this gap.
@@ V-D Derivation of Cooperative Maneuver
VI Simulative Experiments
We integrated the proposed cooperative planning approaches into a new simulation and evaluation framework for fully automated and mixed traffic. The cooperative planners are evaluated and assessed against two baselines. Our primary baseline are non-cooperative (NC) CAVs that leverage shared perception data, but no routing information and adhere to priority rules. This baseline employs the same underlying trajectory planning as is used for the cooperative approaches. Thus, it constitutes a rather strong baseline, which facilitates a sharp evaluation of the cooperative aspect in isolation. In addition, we present the metrics for pure HDV traffic, which is simulated using a learned behavior model. As such, it resembles human driving behavior more closely and exhibits a human-typical agile behavior.
VI-A Simulation Setup
The simulation framework that we used for evaluation of the planning methods builds upon the DeepSIL simulator introduced in [40]. It is based on the ROS2 software distribution [41] and can simulate multiple HDVs as well as CAVs. The simulation has a base time step of .
@@ VI-A Simulation Setup
Connected automated vehicles are simulated using a single-track model with friction and dead time, resembling one of our real-world CAVs. The inputs to this model are calculated by the trajectory planner from [42], which is part of this CAV’s software and therefore provides a realistic CAV behavior. Like HDVs, CAVs do not know the route of other vehicles, even of other CAVs, and assume the most conflicting route. However, the simulated CAVs do not need to approach the intersection slowly due to occlusion, because they receive the intersection EM as in [1].
VI-B Scenario Definition
We define all simulated scenarios at an exemplary intersection in Ulm-Lehr, Germany, as shown in Fig. 6, where also our pilot site (cf. Sec. VII-A) is located. It is a T-junction with a bending main road and a side road. The speed limit is () on the eastern and western part and () on the northern part.
@@ VI-C Quantitative Results
VII Real-World Experiments
@@
VII-A Pilot Site Ulm-Lehr
The real-world experiments were conducted at a pilot site for connected automated driving in Ulm-Lehr, Germany [1]. The intersection is equipped with multiple sensor processing units, each consisting of cameras, radar and lidar sensors as well as a computer for object detection. An edge server resides in the same 5G cellular network cell and a fusion and tracking module [43] combines all object detections into the infrastructure EM, which is provided to the connected vehicles and the cooperative planning modules. The latter also run on the edge server. We employed three CAVs, of which one used the trajectory planning from [42] as mentioned before, and the other two implemented the concept from [44].
The scenario that was investigated in real-world experiments comprises the three CAVs approaching the intersection from each of its accesses, as depicted in Fig. 1. Vehicle approaches the intersection on the main road coming from north and following the road. According to ordinary priority rules, this vehicle has the right of way. Vehicle starts at the other main road access (east) and intents to turn left at the intersection. This vehicle vehicle would have to yield to the oncoming traffic but not to any vehicle coming from the side road. With vehicle approaching the intersection from the west on the minor road, it would have to yield to all cross traffic. As it intents to turn right at the intersection, there is no conflict with but only with . Figure 1 illustrates a typical cooperative maneuver. If vehicle waives its right of way and possibly slows down while approaching, it allows both other CAVs to cross the intersection virtually simultaneously. Depending on the timing, vehicles and can simply traverse the intersection unimpeded. Ideally, vehicle should not be required to stop, but slow down smoothly and accelerate again once the intersection is cleared.
VII-B Evaluation in Real-World
@@
VIII Conclusion
In this paper, an optimization-based and an RL-based approach for cooperative maneuver planning in mixed traffic at urban intersections have been compared. The proposed approaches have been evaluated in simulation and in real-world experiments using three prototype CAVs. Compared to HDVs adhering to priority rules and non-cooperative CAVs, the cooperative maneuvers lead to improved traffic efficiency. In simulated fully automated traffic, a significant reduction of delay due to interaction and of the share of stopped vehicles has been shown. In mixed traffic with CAV penetration, the proposed planners still found a viable cooperative maneuver in about one third of the scenarios. Both cooperative planners demonstrated their real-world applicability in public traffic and yielded a significant reduction in delay due to interaction. Future work will include the evaluation of further cooperative use cases in urban traffic involving human-driven CVs and vulnerable road users.

### v2

@@ 5.3 Reward Function Engineering
with describing the set of reward components and the corresponding weights. The only positive reward component and thus the main driver for learning a nontrivial solution is the velocity reward, which is proportional to the average driving speed clipped to the lane speed limit. A trivial solution that consequently avoids any collision is to stop all vehicles in the scene. This is not a valid solution for cooperative maneuver planning and thus prevented through the idle penalty. In mixed traffic, the training tends to converge to a very conservative behavior, which manifests in all CAVs being stopped far away from the intersection to let HDVs pass. This issue is addressed by the reluctance penalty, which is triggered for CAVs that drive very slowly without any lead vehicle. In order to facilitate learning of reasonable safety distances, the proximity reward penalizes situations in which CAVs get dangerously close. Ultimately, two vehicles colliding triggers the collision penalty and ends the training episode to avoid further positive rewards from being accumulated. The reward weights used throughout this study, at a policy frequency of , are given in Table 2. Details on the definition and tuning of the various reward components can be found in [24, 3].
@@
5.4 Derivation of Cooperative Maneuver
The maneuver representation introduced in Sec. 3 requires the cooperative maneuver to be specified over a sufficient horizon to facilitate local trajectory planning on the CAVs. This requirement for advance planning contrasts with the iterative nature of the RL policy, which requires immediate state feedback. In the present work, we build upon our approach from [26] to bridge this gap.
@@ 5.4 Derivation of Cooperative Maneuver
6 SIMULATIVE EXPERIMENTS
We integrated the proposed cooperative planning approaches into an extended simulation and evaluation framework for fully automated and mixed traffic. The cooperative planners are evaluated and assessed against two baselines. Our primary baseline are non-cooperative (NC) CAVs that leverage shared perception data, but no routing information and adhere to priority rules. This baseline employs the same underlying trajectory planning as is used for the cooperative approaches. Thus, it constitutes a rather strong baseline, which facilitates a sharp evaluation of the cooperative aspect in isolation. In addition, we present the metrics for pure HDV traffic, which is simulated using a learned behavior model. As such, it resembles human driving behavior more closely and exhibits a human-typical agile behavior.
6.1 Simulation Setup
The simulation framework that we used for evaluation of the planning methods builds upon the DeepSIL simulator introduced in [4]. It is based on the ROS2 software distribution [40] and can simulate HDVs as well as CAVs. We extended the framework to support multiple CAVs and cooperative maneuvers. The simulation has a base time step of and runs in real-time, meaning that processing delays from the cooperative maneuver planners directly affect the results. When simulating more than four CAVs, the simulation is slowed down slightly to still be able to run all vehicle trajectory planners in parallel.
@@ 6.1 Simulation Setup
Connected automated vehicles are simulated using a single-track model with friction and dead time, resembling one of our real-world CAVs. The inputs to this model are calculated by the trajectory planner and controller from [5], which is part of this CAV’s software and therefore provides a realistic CAV behavior. Like HDVs, CAVs do not know the route of other vehicles, even of other CAVs, and assume the most conflicting route. However, the simulated CAVs do not need to approach the intersection slowly due to occlusion, because they receive the intersection EM as in [1].
6.2 Scenario Definition
We define all simulated scenarios at an exemplary intersection in Ulm-Lehr, Germany, as shown in Fig. 6, where also our pilot site (cf. Sec. 7.1) is located. It is a T-junction with a bending main road and a side road. The speed limit is () on the eastern and western part and () on the northern part.
@@ 6.3 Quantitative Results
7 REAL-WORLD EXPERIMENTS
@@
7.1 Pilot Site Ulm-Lehr
The real-world experiments were conducted at a pilot site for connected automated driving in Ulm-Lehr, Germany [1]. An aerial view of the site is shown in Fig. 9. The unsignalized intersection is equipped with multiple sensor processing units, each consisting of cameras, radar and lidar sensors as well as a computer for object detection. An edge server resides in the same 5G network cell and a fusion and tracking module [42] combines all object detections into the infrastructure EM, which is provided to the connected vehicles and the cooperative planning modules. The latter also run on the edge server. We employed three CAVs, of which one used the trajectory planning from [5] and the other two implemented the concept from [43].
The scenario that was investigated in real-world experiments comprises the three CAVs approaching the intersection from each of its accesses, as depicted in Fig. 1. Vehicle approaches the intersection on the main road coming from north and following the road. According to ordinary priority rules, this vehicle has the right of way. Vehicle starts at the other main road access (east) and intends to turn left at the intersection. This vehicle would have to yield to the oncoming traffic but not to any vehicle coming from the side road. With vehicle approaching the intersection from the west on the minor road, it would have to yield to all cross traffic. As it intends to turn right at the intersection, there is no conflict with but only with . Figure 1 illustrates a typical cooperative maneuver. If vehicle waives its right of way and possibly slows down while approaching, it allows both other CAVs to cross the intersection virtually simultaneously. Depending on the timing, vehicles and can simply traverse the intersection unimpeded. Ideally, vehicle should not be required to stop, but slow down smoothly and accelerate again once the intersection is cleared.
7.2 Evaluation in Real-World
@@
8 CONCLUSION
In this paper, an optimization-based and an RL-based approach for cooperative maneuver planning in mixed traffic at urban intersections have been evaluated. The proposed approaches have been compared in a realistic simulation and in real-world experiments using three prototype CAVs. Compared to HDVs adhering to priority rules and non-cooperative CAVs, the cooperative maneuvers lead to improved traffic efficiency. In simulated fully automated traffic, a significant reduction of delay due to interaction and of the share of stopped vehicles has been shown. Even in mixed traffic with only CAV penetration, the proposed planners still found a viable cooperative maneuver in about one third of the scenarios, despite not relying on any cooperation or turn direction information from the HDVs. Both cooperative planners demonstrated their real-world applicability in public traffic and yielded a significant reduction in delay due to interaction. Future work will include the evaluation of further cooperative use cases in urban traffic involving human-driven connected vehicles and vulnerable road users.